OpenStudy (anonymous):

x^2-1/4x (needs to be perfect a

OpenStudy (anonymous):

x^2-1/4x (needs to be a perfect square trinomial. help please

Directrix (directrix):

x² - 1/4x = Take half the coefficient of the linear x term and square it. (1/2) * (-1/4)² = ? --> @etorres4

Directrix (directrix):

@etorres4 If you will multiply that, we can wrap up this problem.

Directrix (directrix):

(1/2) * (-1/4)² = ? (-1/4)* (-1/4) * (1/4) = ?

OpenStudy (anonymous):

(x-1/8)^2 -1/64 if you simplify ii that would give the same answer